# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### v5

@@ 4.2.2 Feature Extraction
In Table 4, we delve into an ablation study to examine the impact of different backbones and pre-training methods. Initially, we evaluate different backbones without any pre-training. The backbone of FadNet++ [46], despite its low parameter count of 1.77M, results in an EPE of 0.8633, indicating less precision in disparity estimation. Employing MobilenetV2 100 [51] as a backbone improves the EPE to 0.8227, with a slight increase in computational demand (31.83G flops) and parameters (2.78M). The backbone of PSMNet [5], while the most computationally intensive (130.87G flops) with the highest number of parameters (4.46M), achieves a further reduction in EPE to 0.8068, suggesting that the complexity of the network can contribute to the enhancement of stereo matching performance. Furthermore, we delve into the impact of pre-training on the MobilenetV2 100 backbone. Without pre-training, the EPE remains at 0.8227. However, when pre-trained, we observe a substantial improvement, as the EPE drops to 0.7132. This underscores the value of leveraging large-scale image datasets to enhance feature extraction capabilities before fine-tuning specific tasks like stereo matching. In conclusion, our ablation study highlights the critical role of both the architecture’s complexity and pre-training in refining the stereo matching model’s accuracy.
Table 5: Ablation study results of cost volume on the finalpass of SceneFlow test datasets [28]. For these experiments, one-quarter features from stereo images are utilized to construct the cost volume. Gwc represents Group-wise correlation volume [13]. Cat stands for Concatenation volume [5]. G8-C16, G16-C24 and G32-C48 combine Group-wise correlation volume and Concatenation volume [13]. Channel represents the channel of cost volume. Dims refers to the dimensions of cost volume.
@@
Difference 3D - 17.44G 2.40M 1.02
Correlation 3D - 24.80G 4.01M 0.81
Interlaced8 4D 8 126.87G 2.83M 0.70
Gwc8 4D 8 31.83G 2.78M 0.71
Gwc16 4D 16 75.28G 3.89M 0.66
Gwc24 4D 24 147.53G 5.75M 0.63
Gwc32 4D 32 248.60G 8.34M 0.62
Gwc48 4D 48 537.15G 15.73M 0.60
Cat24 4D 24 148.36G 5.78M 0.65
Cat48 4D 48 537.99G 15.76M 0.61
Cat64 4D 64 941.79G 26.09M 0.60
G8-C16 4D 24 148.36G 5.78M 0.62
G16-C24 4D 40 379.30G 11.69M 0.60
G32-C48 4D 80 1460.80G 39.37M 0.60
@@ 4.2.3 Cost Construction
In Table 5, an ablation study on various cost volume strategies for stereo matching is presented. The study begins with simpler 3D cost volume methods like Difference [21] and Correlation [46], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs. This suggests that while efficient, these methods may lack the nuanced disparity capture necessary for complex scenes. The Interlaced8 model, introduced by MobileStereoNet [39], achieves the same EPE comparable to the Gwc8 model. However, its computational expense is substantially higher, with a flop count of 126.87G which is significantly larger than that of the Gwc8 model. The group-wise correlation and concatenation models demonstrate a clear trend: as the channel depth increases, the EPE improves, indicating improved disparity estimations through richer feature capture. The combined volume (G8-C16) offers a more optimal balance between computational load and disparity estimation accuracy, which achieves an EPE of 0.62. G16-C24 and G32-C48 do not result in significant improvement in EPE, which stabilizes at 0.60, despite a dramatic increase in computational load, especially for G32-C48 which demands 1460.80 Gflops and has 39.37M parameters. These results highlight the delicate balance between accuracy and computational efficiency in designing cost volumes for disparity estimation. While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
Table 6: Ablation study results of disparity regression and refinement on the finalpass of SceneFlow test datasets [28]. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [42, 25, 51].
@@
ArgMin None 26.37G 2.69M 0.76
ArgMin RGBRefine [52] 53.15G 2.81M 0.72
ArgMin Context 31.83G 2.78M 0.71
ArgMin Context+RGBRefine [52] 58.61G 2.89M 0.69
ArgMin Context+DRNetRefine [52] 58.57G 2.89M 0.69
ArgMin ConvGRU [25, 51] 1363.71G 12.51M 0.46
@@ 4.2.4 Disparity Regression and Refinement
The Differentiable ArgMin [20, 5, 13, 40, 50, 12, 58, 39, 46] introduced by GCNet [20], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. As shown in Table 6, in this ablation study on disparity refinement for the SceneFlow test datasets, various strategies show differing impacts on model performance. Without refinement, the model has an EPE of 0.76. RGBRefine and Context methods slightly improve EPE to 0.72 and 0.71, respectively, with a modest increase in computational resources. Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration. However, employing ConvGRU refinement leads to a substantial improvement, reducing EPE to 0.46, albeit at a significant cost in computational complexity (1363.71 Gflops) and model size (12.51M). This highlights a trade-off between accuracy improvements and increased computational demands.
@@ 4.3.1 Necessity of Comprehensive Ablation Study
In the evolving landscape of deep stereo matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies. A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios. For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately. Similarly, the impact of varying backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound. Understanding the specific contributions of each component is crucial for building more efficient and effective stereo matching systems.
@@ 4.3.2 Necessity of A Strong Baseline
A strong baseline in deep stereo matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses. In summary, a strong baseline is essential for meaningful advancements in deep stereo matching, ensuring that new developments are substantial, accurately assessed, and broadly applicable.
@@ 5 A Strong Pipeline: StereoBase
In light of our comprehensive analysis, the goal of this section is to establish a strong baseline model that surpasses existing standards in performance. StereoBase, our proposed solution, embodies this objective.
@@ 5.1 Pipeline
We build a strong baseline: StereoBase. StereoBase begins with a carefully curated data preprocessing step. For SceneFlow, KITTI2012, and KITTI2015, we implement an optimized mix of data augmentation techniques, including RandomCrop, ColorAug, and Erase Transform. This step ensures that the input data is not only varied and representative of real-world scenarios but also tailored to enhance stereo matching performance. Given the left and the right images, the pre-trained MobileNetV2 [48] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume. The cost volume utilizes a combined volume, which combines group-wise correlation volume with an 8-channel and concatenation volume with 16 channels, to ensure an optimal balance between computational load and disparity estimation accuracy. Hourglass networks [51] were implemented for cost aggregation, while convGRU [51] strategies were applied for the final disparity regression.
Table 7: Results on SceneFlow [28], KITTI 2012 [11], KITTI 2015 [29] leaderboard, and DrivingStereo [55]. All results on DrivingStereo [55] are derived using the OpenStereo. Underline refers to evaluation in the non-occluded regions only. Bold: Best.
@@
5.2 Comparison with State-of-the-art
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow, KITTI2012, 2015, and DrivingStereo. On the SceneFlow test set, we achieve a new SOTA EPE of 0.34. The quantitative comparisons, as summarized in Table 7, clearly illustrate the edge of StereoBase in handling complex stereo matching scenarios with greater precision. Further, we submitted our results to the KITTI2012 and 2015 leaderboard, where StereoBase outperformed all published methods across all metrics. On KITTI2015, our StereoBase outperforms IGEV [51] by 9.43% on D1-all metric, respectively. In addition, we evaluate the generalization performance of StereoBase. As shown in Table 8, StereoBase exhibited exceptional performance in a zero-shot setting, achieving state-of-the-art results. This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
@@ 6 Conclusion
This paper introduces OpenStereo, a codebase designed for deep stereo matching. Our initial endeavor involved the re-implementation of the most state-of-the-art methods within the OpenStereo framework. This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo matching methodologies. Drawing on the insights gained from our exhaustive ablation studies, we proposed StereoBase. StereoBase not only demonstrates the capabilities of our platform but also sets a new standard in the field for future research and development. Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo matching community, fostering innovation and facilitating more effective and efficient research.

### v6

@@ 4.2.2 Feature Extraction
As shown in Table 2, pretraining the backbone is crucial for stereo matching as it enhances the model’s ability to extract robust and informative features. Furthermore, the choice of backbone significantly influences the model’s performance and computational efficiency. MobilenetV2 [42] and EfficientNetV2 [48] are lightweight CNNs that are particularly efficient in extracting local features, which are crucial for stereo matching. Their designs allow them to perform well with relatively low computational complexity. RepViT [50] is a Transformer-based architecture, which excels in capturing long-range dependencies and global context. While RepViT captures global features well, it might struggle with the fine-grained, pixel-level accuracy required for precise stereo matching. MPViT [27] combines the strengths of both CNNs and Transformers. The CNN components effectively capture local features, while the Transformer components excel in modeling global context. This hybrid approach allows MPViT to leverage the advantages of both architectures, resulting in the lowest EPE. In summary, MobilenetV2 [42] offers a good balance for applications with limited computational resources, while more complex architectures like EfficientNetV2 and MPViT provide superior accuracy at the cost of higher computational requirements. To the best of our knowledge, our work is the first to explore the transformer-based feature extraction and the combination of CNN and transformer feature extraction for stereo matching.
Table 3: Ablation study of cost volume on the finalpass of SceneFlow test datasets [35]. Gwc represents Group-wise correlation volume [18]. Cat stands for Concatenation volume [6]. G8-C16, G16-C24, and G32-C48 combine Gwc volume and Cat volume [18]. Channel and Dims represent the channel and dimensions of the cost volume, respectively. FLOPs are calculated at a resolution of $544\times 960544 × 960$.
@@
Difference 3D - 38.68G 2.40M 1.02
Correlation 3D - 54.99G 4.01M 0.81
Interlaced8 4D 8 288.52G 2.83M 0.70
Gwc8 4D 8 70.58G 2.78M 0.72
Gwc16 4D 16 166.92G 3.89M 0.66
Gwc24 4D 24 327.13G 5.75M 0.63
Gwc32 4D 32 551.23G 8.34M 0.62
Gwc48 4D 48 1191.07G 15.73M 0.60
Cat24 4D 24 328.97G 5.78M 0.65
Cat48 4D 48 1192.93G 15.76M 0.61
Cat64 4D 64 2088.31G 26.09M 0.60
G8-C16 4D 24 328.96G 5.78M 0.62
G16-C24 4D 40 841.05G 11.69M 0.60
G32-C48 4D 80 3239.17G 39.37M 0.60
@@ 4.2.3 Cost Construction
In Table 3, an ablation study on various cost volume strategies for stereo matching is presented. For these experiments, one-quarter of stereo image features are used to construct the cost volume. The study begins with simpler 3D cost volume methods: Difference [26] and Correlation [52], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs. This suggests that while efficient, these methods may lack the nuanced disparity capture necessary for complex scenes. The Interlaced8 model, introduced by MobileStereoNet [46], achieves the same EPE comparable to the Gwc8 model. However, its computational expense is substantially higher, with a flop count of 288.52G, significantly larger than that of the Gwc8 model. The group-wise correlation and concatenation models demonstrate a clear trend: as the channel depth increases, the EPE improves, indicating improved disparity estimations through richer feature capture. The combined volume (G8-C16) offers a more optimal balance between computational load and disparity estimation accuracy, which achieves an EPE of 0.62. G16-C24 and G32-C48 do not significantly improve EPE, despite a dramatic increase in computational load, especially for G32-C48, which demands 3239.17Gflops and has 39.37M parameters. These results highlight the delicate balance between accuracy and computational efficiency in designing cost volumes for disparity estimation. While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
Table 4: Ablation study of disparity regression and refinement on the finalpass of SceneFlow test datasets [35]. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [33, 60].
@@
ArgMin None 58.47G 2.69M 0.76
ArgMin RGBRefine [61] 117.85G 2.81M 0.72
ArgMin Context 70.58G 2.78M 0.71
ArgMin Context+RGBRefine [61] 129.95G 2.89M 0.69
ArgMin Context+DRNetRefine [61] 129.88G 2.89M 0.69
ArgMin ConvGRU [33, 60] 3023.88G 12.51M 0.46
@@ 4.2.4 Disparity Regression and Refinement
The Differentiable ArgMin [25, 6, 18, 47, 59, 15, 70, 46, 52] introduced by GCNet [25], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. As shown in Table 4, various strategies show differing impacts on model performance in this ablation study on disparity refinement for the SceneFlow test datasets. Without refinement, the model has an EPE of 0.76. RGBRefine and Context methods slightly improve EPE to 0.72 and 0.71, respectively, with a modest increase in computational resources. Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration. However, employing ConvGRU refinement substantially improves EPE 0.46, albeit at a significant cost in computational complexity (3023.88 Gflops) and model size (12.51M). This highlights a trade-off between accuracy improvements and increased computational demands.
@@ 4.3.1 Necessity of Comprehensive Ablation Study
In the evolving landscape of deep stereo matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies. A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios. For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately. Similarly, the impact of varying backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound. Understanding the specific contributions of each component is crucial for building more efficient and effective stereo-matching systems.
@@ 4.3.2 Necessity of A Strong Baseline
A strong baseline in deep stereo-matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses. In summary, a strong baseline is essential for meaningful advancements in deep stereo matching, ensuring that new developments are substantial, accurately assessed, and broadly applicable.
@@ 5.1 Pipeline
In light of our comprehensive analysis, the goal of this section is to establish a strong baseline model that surpasses existing standards in performance. StereoBase embodies this objective. As shown in Figure 5, given the left and the right images, the pre-trained MobileNetV2 [57] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume. The G8-C16 cost volume is utilized to achieve an optimal balance between computational load and disparity estimation accuracy. Hourglass networks [60] were implemented for cost aggregation, while convGRU [60] strategies were applied for the final disparity regression.
Figure 5: Overview of our proposed StereoBase. GwcVolume represents Group-wise correlation volume [18]. CGEV refers to Combined Geometry Encoding Volume [60].
Table 5: Results on SceneFlow [35], KITTI 2012 [14], KITTI 2015 [36] leaderboard, and DrivingStereo [66]. All results on DrivingStereo [66] are derived using the OpenStereo. Underline refers to evaluation in the non-occluded regions only STTR [31]. Bold: Best.
@@
GANet+ADL [63] 0.50 0.98 1.29 1.38 2.38 1.55 - -
NMRF-Stereo [17] 0.45 1.01 1.35 1.28 3.13 1.59 - -
Selective-IGEV [53] 0.44 1.07 1.38 1.33 2.61 1.55 - -
MoCha-Stereo [8] 0.41 1.06 1.35 1.36 2.43 1.53 - -
@@
5.2 Comparison with State-of-the-art Methods
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow [35], KITTI2012 [14], 2015 [36], and DrivingStereo [66]. On the SceneFlow [35] test set, we achieve a new SOTA EPE of 0.34. The quantitative comparisons, as summarized in Table 5, clearly illustrate the edge of StereoBase in handling complex stereo-matching scenarios with greater precision. Further, we submitted our results to the KITTI2012 [14] and 2015 [36] leaderboards, where StereoBase outperformed all published methods across all metrics. On KITTI2015 [36], our StereoBase outperforms IGEV [60] by 9.43% on D1-all metric, respectively. In addition, we evaluate the generalization performance of StereoBase. As shown in Table 6, StereoBase exhibited exceptional performance in a zero-shot setting. This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
@@ 6 Conclusion
This paper introduces OpenStereo, a benchmark designed for deep stereo matching. Our initial endeavor involved re-implementing the most state-of-the-art methods within the OpenStereo framework. This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo-matching methodologies. Drawing on the insights gained from our exhaustive ablation studies, we proposed StereoBase. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and performs best across all metrics. In addition, StereoBase has strong cross-dataset generalization. StereoBase not only demonstrates the capabilities of our platform but also sets a new standard in the field for future research and development. Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo-matching community, fostering innovation and facilitating more effective and efficient research.
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems.